• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name A. Deur

    Log entry time 00:06:06 on October 4,2002

    Entry number 87293

    keyword=Shift Summary

    Helium level instable for Q1 left HRS. Ed is working on it.

    ~19:00 MCC called saying they will be able to deliver the beam in about 1 hour. But Q1 problem is not expected to be solved by then.

    22:45 Sweep of Hall A (Q1 problem is still not understood)
    23:41 Power Permit
    Move target from L2 bottom cell to thick carbon but the movement is not working. We called the target on call.
    We will wait Eugene for a Moller Polarimetry